# Millimeter emission from supermassive black hole coronae

### v1

@@ B.5 NGC 985
This galaxy located at $D\approx 190$ Mpc hosts a SMBH with $M_{\mathrm{BH}}=2.2\times 10^{8}$ M (Inoue & Doi 2018, and references therein). Inoue & Doi (2018) reported an excess in the SED of this source at $\sim$100 GHz from which they inferred the presence of a corona. The dust component is essentially unconstrained due to a very bright FIR flux from the whole galaxy (Ichikawa et al. 2019). The high-resolution ALMA data favours a corona with a harder electron population, consistent with $p\approx 2.1$ (Inoue & Doi 2018). Trying to fit this parameter hits the lower limit in the priors, and thus we fix it to $p=2.1$. Following Inoue & Doi (2018), we also fix $kT_{\mathrm{c}}=29$ keV and $\tau_{\mathrm{T}}=3.5$. The inferred properties of the corona are: $r_{\mathrm{c}}=116^{+44}_{-18}$ ($R_{\mathrm{c}}\approx 1.5$ ld) and $\log{\delta}=-2.88^{+0.15}_{-0.25}$, leading to $B\sim 13.4$ G and $\sigma\sim 0.013$. The corona SED component peaks at $\nu_{\mathrm{p}}=94\pm 15$ GHz, with a peak flux density of $S_{\mathrm{p}}=1.4\pm 0.2$ mJy, and has a bolometric synchrotron luminosity of $L_{\mathrm{c}}\approx 3\times 10^{40}$ erg s-1. We estimate a hard X-ray luminosity $L_{\mathrm{14-150\,keV}}\approx 4\times 10^{45}$ erg s-1, corresponding to an Eddington ratio of $\lambda_{\mathrm{Edd}}\sim 0.011$. The cornerplot of the fit is shown in Fig. 15 and the data used is listed in Table 6.
Table 6: Flux densities and sizes of the flux extraction regions for NGC 985. The Type column indicates if the values were used for the fit, considered as upper limits (UL) or lower limits (LL). References: (a)this work; (b)Kawamuro et al. (2022); (c)Ichikawa et al. (2019).
@@ B.6 NGC 1068
This galaxy is located at $D\approx 14.6$ Mpc ($z=0.0033$) and host a SMBH with $M_{\mathrm{BH}}=(1.66\pm 0.01)\times 10^{7}$ M (Gallimore et al. 2024). This source has been the subject of very extensive research, given that it was the first RQ galaxy to be identified as a neutrino source (IceCube Collaboration et al. 2022). It has also been extensively studied in X-rays, revealing the presence of a corona (e.g. Bauer et al. 2015). Furthermore, Inoue et al. (2020) modelled the radio continuum data and concluded that the synchrotron emission from the corona was responsible for the sub-mm emission detected by ALMA, which means that the corona peaks at a rather high frequency in this source compared to others. In a more recent work, Mutie+2025 presented a consistent SED of the source matched in resolution and uv-coverage, confirming the presence of corona emission. For completeness, we reproduce the fluxes from their work in Table 7.
The nuclear region of this source is particularly complex, and so is its SED at low frequencies (Mutie et al. 2025). For this reason, here we consider only high-resolution observations at $\nu\geq 15$ GHz. Due to the lack of FIR data to constrain the dust component, we fix $\beta=1.78$ and $\nu_{\tau_{1}}=800$ GHz as with the other sources. We refer to García-Burillo et al. (2016) for a discussion of the dust emission, although we caution that these authors did not take into account the coronal emission in their analysis. In addition, we fix $\alpha_{\mathrm{sy}}=-0.5$ as otherwise the spectral index hits the hard limit set on $-0.5$. We obtain $r_{\mathrm{c}}=67.5\pm 5.5$ ($R_{\mathrm{c}}\approx 0.065$ ld) and $\log{\delta}=-0.98\pm 0.13$, leading to $B\sim 158$ G and $\sigma\sim 1.1$. The corona SED peaks at $\nu_{\mathrm{p}}\simeq 568\pm 55$ GHz, with a peak flux density of $S_{\mathrm{p}}=10.1\pm 1.9$ mJy, and has a bolometric synchrotron luminosity of $L_{\mathrm{c}}\approx 4.4\times 10^{39}$ erg s-1. Given that the peak frequency is much greater than 100 GHz, we estimate the hard X-ray luminosity using the relation from Kawamuro et al. (2022); we obtain a luminosity $L_{\mathrm{14-150\,keV}}\approx 2.3\times 10^{43}$ erg s-1, corresponding to an Eddington ratio of $\lambda_{\mathrm{Edd}}\sim 0.011$. The cornerplot of the fit is shown in Fig. 16.
Table 7: Flux densities and sizes of the flux extraction regions for NGC 1068. The Type column indicates if the values were used for the fit, considered as upper limits (UL) or lower limits (LL). References: (a)Mutie et al. (2025).
@@ B.7 NGC 3227
This galaxy located at $D\approx 23$ Mpc (Ricci et al. 2023) hosts a $M_{\mathrm{BH}}=4.2\times 10^{7}$ M SMBH (Behar et al. 2015). Behar et al. (2015) showed that this source has a significant excess at $\sim$100 GHz. Sani et al. (2012) reported its flux density at 3 mm as observed by the Plateau de Bure with a $\sim$1” resolution. In addition, flux densities from ALMA observations at B6 and B7 were reported in Alonso-Herrero et al. (2019). According to the very high angular resolution observations at B3 in Ricci et al. (2023), the flux from the core region is $\sim$0.85 mJy. The dust component is poorly constrained due to a very bright FIR flux from the whole galaxy (Ciesla et al. 2012). However, Alonso-Herrero et al. (2019) fitted the IR SED with a dust model that has a flux of $\sim$10 mJy at $10^{3}$ GHz (their Fig. 4), so that we can at least check that our model fitting is consistent with that derived independently using dust modelling of mid-IR low-resolution observations. Unfortunately, it is not possible to fit a dataset with very high angular resolution ($<0.1"$) due to the limited number of data points. We thus use as reference points with resolutions $\sim$0.4”–1”, which is not optimal for detecting the relatively faint corona emission. Thus, the properties of the corona are only loosely constrained as $r_{\mathrm{c}}=67^{+58}_{-31}$ ($R_{\mathrm{c}}\approx 0.06$ ld) and $\log{\delta}=-2.29^{+0.67}_{-0.41}$, leading to $B\sim 36$ G and $\sigma\sim 0.05$. The corona SED component peak is very poorly constrained at $\nu_{\mathrm{p}}=185\pm 91$ GHz, with a peak flux density of $S_{\mathrm{p}}=0.8\pm 0.5$ mJy, and has a bolometric synchrotron luminosity of $L_{\mathrm{c}}\approx 2\times 10^{38}$ erg s-1. Interestingly, the source is undetected in Q band with the VLA in A array configuration (project code AG568), giving an UL to the compact emission produced at 43.3 GHz, which is consistent with the SED fit of the corona component. We estimate a hard X-ray luminosity $L_{\mathrm{14-150\,keV}}\approx 9\times 10^{42}$ erg s-1, corresponding to an Eddington ratio of $\lambda_{\mathrm{Edd}}\sim 0.004$. The cornerplot of the fit is shown in Fig. 17.
Table 8: Flux densities and sizes of the flux extraction regions for NGC 3227. The Type column indicates if the values were used for the fit, considered as upper limits (UL) or lower limits (LL). References: (a)Behar et al. (2015); (b)Alonso-Herrero et al. (2019); (c)this work; (d)Sani et al. (2012); (e)Bontempi et al. (2012); (f)Smith et al. (2020); (g)Ricci et al. (2023); (h)Ciesla et al. (2012).
@@
8.46 $15.4\pm 0.4$ $2.6\times 2.5$ UL (c)
14.9 $6.5\pm 1.1$ $0.89\times 0.56$ yes (c)
@@
100 $0.85\pm 0.12$ $0.08\times 0.08$ LL (g)
@@ Appendix C Parameter space and cornerplots
We sample a broad range in the parameter space to ensure that all possible solutions are explored. A list of parameters and their allowed ranges is presented in Table 1. We highlight that we fit the logarithm of the parameter for those that can spread over several orders of magnitude (namely the normalisations and $\delta$).

### v2

@@ B.5 NGC 985
This galaxy located at $D\approx 190$ Mpc hosts an SMBH with $M_{\mathrm{BH}}=2.2\times 10^{8}$ M (Inoue & Doi 2018, and references therein). Inoue & Doi (2018) reported an excess in the SED of this source at $\sim$100 GHz from which they inferred the presence of a corona. The dust component is essentially unconstrained due to a very bright FIR flux from the whole galaxy (Ichikawa et al. 2019). The high-resolution ALMA data favours a corona with a harder electron population, consistent with $p\approx 2.1$ (Inoue & Doi 2018). Trying to fit this parameter hits the lower limit in the priors, and thus we fix it to $p=2.1$. Following Inoue & Doi (2018), we also fix $kT_{\mathrm{c}}=29$ keV and $\tau_{\mathrm{T}}=3.5$. The inferred properties of the corona are: $r_{\mathrm{c}}=111^{+33}_{-14}$ ($R_{\mathrm{c}}\approx 1.4$ ld) and $\log{\delta}=-2.03^{+0.13}_{-0.22}$, leading to $B\sim 14.4$ G and $\sigma\sim 0.014$. The corona SED component peaks at $\nu_{\mathrm{p}}=100\pm 15$ GHz, with a peak flux density of $S_{\mathrm{p}}=1.4\pm 0.2$ mJy, and has a bolometric synchrotron luminosity of $L_{\mathrm{c}}\approx 4\times 10^{40}$ erg s-1. We estimate a hard X-ray luminosity $L_{\mathrm{14-150\,keV}}\approx 4\times 10^{45}$ erg s-1, corresponding to an Eddington ratio of $\lambda_{\mathrm{Edd}}\sim 0.011$. The cornerplot of the fit is shown in Fig. 15 and the data used is listed in Table 6.
Table 6: Flux densities and sizes of the flux extraction regions for NGC 985. The Type column indicates if the values were used for the fit as observed (yes), considered as upper limits (UL) or lower limits (LL). References: (a)this work; (b)Inoue & Doi (2018); (c)Ichikawa et al. (2019).
@@ B.6 NGC 1068
This galaxy is located at $D\approx 14.6$ Mpc ($z=0.0033$) and hosts an SMBH with $M_{\mathrm{BH}}=(1.66\pm 0.01)\times 10^{7}$ M (Gallimore et al. 2024). This source has been the subject of very extensive research, given that it was the first RQ galaxy to be identified as a neutrino source (IceCube Collaboration et al. 2022). It has also been extensively studied in X-rays, revealing the presence of a corona (e.g. Bauer et al. 2015). Furthermore, Inoue et al. (2020) modelled the radio continuum data and concluded that the synchrotron emission from the corona was responsible for the sub-mm emission detected by ALMA, which means that the corona peaks at a rather high frequency in this source compared to others. In a more recent work, Mutie et al. (2025) presented a consistent SED of the source matched in resolution and uv-coverage, confirming the presence of corona emission. For completeness, we reproduce the fluxes from their work in Table 7.
The nuclear region of this source is particularly complex, and so is its SED at low frequencies (Mutie et al. 2025). For this reason, here we consider only high-resolution observations at $\nu\geq 15$ GHz. Due to the lack of FIR data to constrain the dust component, we fix $\beta=1.78$ and $\nu_{\tau_{1}}=800$ GHz as with the other sources. We refer to García-Burillo et al. (2016) for a discussion of the dust emission, although we caution that these authors did not take into account the coronal emission in their analysis. In addition, we fix $\alpha_{\mathrm{sy}}=-0.5$ as otherwise the spectral index hits the hard limit set on $-0.5$. We obtain $r_{\mathrm{c}}=67.5\pm 5.5$ ($R_{\mathrm{c}}\approx 0.065$ ld) and $\log{\delta}=-0.98\pm 0.13$, leading to $B\sim 158$ G and $\sigma\sim 1.1$. The corona SED peaks at $\nu_{\mathrm{p}}\simeq 568\pm 55$ GHz, with a peak flux density of $S_{\mathrm{p}}=10.1\pm 1.9$ mJy, and has a bolometric synchrotron luminosity of $L_{\mathrm{c}}\approx 4.4\times 10^{39}$ erg s-1. Given that the peak frequency is much greater than 100 GHz, we estimate the hard X-ray luminosity using the relation from Kawamuro et al. (2022); we obtain a luminosity $L_{\mathrm{14-150\,keV}}\approx 2.3\times 10^{43}$ erg s-1, corresponding to an Eddington ratio of $\lambda_{\mathrm{Edd}}\sim 0.011$. The cornerplot of the fit is shown in Fig. 16. We note that this fit is quite similar to the one obtained by Inoue et al. (2020) and Mutie et al. (2025).
Table 7: Flux densities and sizes of the flux extraction regions for NGC 1068. The Type column indicates if the values were used for the fit as observed (yes), considered as upper limits (UL) or lower limits (LL). References: (a)Mutie et al. (2025).
@@ B.7 NGC 3227
This galaxy located at $D\approx 23$ Mpc (Ricci et al. 2023) hosts a $M_{\mathrm{BH}}=4.2\times 10^{7}$ M SMBH (Behar et al. 2015). Behar et al. (2015) showed that this source has a significant excess at $\sim$100 GHz. Sani et al. (2012) reported its flux density at 3 mm as observed by the Plateau de Bure with a $\sim$1” resolution. In addition, flux densities from ALMA observations at B6 and B7 were reported in Alonso-Herrero et al. (2019). According to the very high angular resolution observations at B3 in Ricci et al. (2023), the flux from the core region is $\sim$0.85 mJy. The dust component is poorly constrained due to a very bright FIR flux from the whole galaxy (Ciesla et al. 2012). However, Alonso-Herrero et al. (2019) fitted the IR SED with a dust model that has a flux of $\sim$10 mJy at $10^{3}$ GHz (their Fig. 4), so that we can at least check that our model fitting is consistent with that derived independently using dust modelling of mid-IR low-resolution observations. Unfortunately, it is not possible to fit a dataset with very high angular resolution ($<0.1"$) due to the limited number of data points. We thus use as reference points with resolutions $\sim$0.4”–1”, which is not optimal for detecting the relatively faint corona emission. Thus, the properties of the corona are only loosely constrained as $r_{\mathrm{c}}=67^{+48}_{-26}$ ($R_{\mathrm{c}}\approx 0.06$ ld) and $\log{\delta}=-2.07^{+0.56}_{-0.48}$, leading to $B\sim 46$ G and $\sigma\sim 0.09$. The corona SED component peak is very poorly constrained at $\nu_{\mathrm{p}}=233\pm 91$ GHz, with a peak flux density of $S_{\mathrm{p}}=0.94\pm 0.48$ mJy, and has a bolometric synchrotron luminosity of $L_{\mathrm{c}}\approx 3\times 10^{38}$ erg s-1. Interestingly, the source is undetected in Q band with the VLA in A array configuration (project code AG568), giving an UL to the compact emission produced at 43.3 GHz, which is consistent with the SED fit of the corona component. We estimate a hard X-ray luminosity $L_{\mathrm{14-150\,keV}}\approx 8\times 10^{42}$ erg s-1, corresponding to an Eddington ratio of $\lambda_{\mathrm{Edd}}\sim 0.004$. The cornerplot of the fit is shown in Fig. 17.
Table 8: Flux densities and sizes of the flux extraction regions for NGC 3227. The Type column indicates if the values were used for the fit as observed (yes), considered as upper limits (UL) or lower limits (LL). References: (a)Behar et al. (2015); (b)Alonso-Herrero et al. (2019); (c)this work; (d)Sani et al. (2012); (e)Bontempi et al. (2012); (f)Smith et al. (2020); (g)Sargent et al. (2025); (h)Ciesla et al. (2012).
@@
6.0 $7.89\pm 0.44$ $0.30\times 0.23$ LL (g)
8.46 $15.4\pm 0.5$ $2.6\times 2.6$ UL VLA(c)
8.5 $12.2\pm 1.3$ $0.2\times 0.2$ LL (a)
10.0 $3.75\pm 0.20$ $0.18\times 0.16$ LL (g)
14.9 $6.5\pm 1.1$ $0.89\times 0.56$ yes VLA(c)
@@
95.2 $2.1\pm 0.7$ $0.66\times 0.45$ UL ALMA(c)
97.5 $1.28\pm 0.35$ $0.13\times 0.09$ LL ALMA(c)
107.9 $1.27\pm 0.25$ $0.16\times 0.11$ LL ALMA(c)
@@ Appendix C Parameter space and cornerplots
We sample a broad range in the parameter space to ensure that all possible solutions are explored. A list of parameters and their allowed ranges is presented in Table 1. We highlight that we fit the logarithm of those parameter that can spread over several orders of magnitude (namely the normalisations and $\delta$).